OpenStudy (anonymous):

If u=(2,1,1) and v=(4,k,2), are orthogonal, find the value of k.

OpenStudy (anonymous):

2*4+1*K+1*2=0 8+K+2=0 K+10=0 K=-10 as the a1a2+b1b2+c1c2=0

OpenStudy (anonymous):

can u please tell me is thevalue of k given in u rbuk?is dat matches wid dis calulated one?